OpenStudy (anonymous):

Fifteen is % of 120.

OpenStudy (anonymous):

12.5

OpenStudy (studygurl14):

15 = WP(120) (WP = what percent) solve for WP

OpenStudy (anonymous):

12.5

OpenStudy (studygurl14):

@sethRB26DETT @Dubba2000 do not give direct answers

OpenStudy (texaschic101):

any kind of percent problem can be solved by using the percent formula is/of = %/100 15 is what % of 120 is = 15 of = 120 % = x now we sub 15/120 = x/100 -- this is a proportion, so we will cross multiply 120x = 1500 x = 1500/120 x = 12.5 its that simple. Like I said, any kind of percent problem can be solved with this formula...just make sure you put the right numbers in the right place